If Bounder Dog Supplies, Inc. purchased inventory at $2,000 list price and the terms were 4/10, n/60, what would be the value associated with the inventory if payment was made within 10 days? a. 1920 b. 1947 c. 2080 d. 2000
Answer #1

List price = 2000

Terms is 4/10 net 60 which means that 4% discount will be given if payment is made within 10 days and total credit period allowed is 60 days.

In the given question, payment is made on 10th day, So, 4% discount will be provided.

Discount on inventory = List price * Discount rate

= 2000*4% = 80

Value of inventory = List price - Discount price

= 2000 - 80 = 1920

Answer is a. 1920